# The parametric complexity of operator learning

### v3

@@ 5.1. HJ-Net Beats The Curse of Parametric Complexity
The error and complexity estimate of Theorem 5.1 implies that for moderate dimensions $d$, and for sufficiently smooth input functions $u_{0}\in\mathcal{F}\subset C^{r}_{\mathrm{per}}$, with $r>3d+1$, the overall complexity of this approach scales at most linearly in $\epsilon^{-1}$: Indeed, the mapping of the data points $(q^{j}_{0},p^{j}_{0},z^{j}_{0})\mapsto(q^{j}_{t},p^{j}_{t},z^{j}_{t})$ for $j=1,\dots,N$ requires $N$ forward-passes of the neural network $\Psi_{t}$, which has $O(\epsilon^{-(2d+1)/r}\log(\epsilon^{-1}))$ internal degrees of freedom. Since $N=O(\epsilon^{-d/r})$ the computational complexity of this mapping is thus bounded by $O(\epsilon^{-(3d+1)/r}\log(\epsilon^{-1}))=O(\epsilon^{-1})$. A naive implementation of the pruning algorithm requires at most $O(N^{2})=O(\epsilon^{-2d/r})=O(\epsilon^{-1})$ comparisons. The computational complexity of the reconstruction by the moving least squares method with $m\leq N$ (pruned) interpolation points and with $M\sim\epsilon^{-d/r}$ evaluation points can be bounded by $O(m+M)=O(\epsilon^{-d/r}+M)=O(\epsilon^{-1})$ [53, last paragraph of Chapter 4.2]. In particular, the overall complexity to obtain an $\epsilon$-approximation of the output function $\mathcal{S}_{t}(u_{0})\approx\mathcal{S}_{t}^{\dagger}(u_{0})$ at e.g. the encoding points $Q_{\epsilon}$ (with $M=|Q_{\epsilon}|\sim\epsilon^{-d/r}$) is at most linear in $\epsilon^{-1}$.
@@ 5.3. Moving Least Squares Reconstruction Error
the pruned moving least squares approximant $f_{{z},{Q}}$ of Algorithm 2 with interpolation data points $Q=\{q^{1},\dots,q^{N}\}$ and data values $z=\{z^{1},\dots,z^{N}\}$, satisfies
@@ 6. Conclusions
The first contribution of this work is to study the curse of dimensionality in the context of operator learning, here interpreted in terms of the infinite dimensional nature of the input space. We state a theorem which, for the first time, establishes a general form of the curse of parametric complexity, a natural scaling limit of the curse of dimensionality in high-dimensional approximation, characterized by lower bounds which are exponential in the desired error. The theorem demonstrates that in general it is not possible to obtain complexity estimates, for the size of the approximating neural operator, that grow algebraically with inverse error unless specific structure in the underlying operator is leveraged; in particular, we prove that this additional structure has to go beyond $C^{r}$- or Lipschitz-regularity. This considerably generalizes and strengthens earlier work on the curse of parametric complexity in [32], where a mild form of this curse had been identified for PCA-Net. As shown in the present work, our result applies to many proposed operator learning architectures including PCA-Net, the FNO, and DeepONet and recent nonlinear extensions thereof. The lower complexity bound in this work is obtained for neural operator architectures based on standard feedforward ReLU neural networks, and could likely be extended to feedforward architectures employing piecewise polynomial activations. It is of note that algebraic complexity bounds, which seemingly overcome the curse of parametric complexity of the present work, have recently been derived for the approximation of Lipschitz operators [49]; these results build on non-standard neural network architectures with either superexpressive activation functions, or non-standard connectivity, and therefore do not contradict our results.
@@
The second contribution of this paper is to present an operator learning framework for Hamilton-Jacobi equations, and to provide a complexity analysis demonstrating that the methodology is able to tame the curse of parametric complexity for these PDE. We present the ideas in a simple setting, and there are numerous avenues for future investigation. For example, as pointed out in Remark 3.5, one main limitation of the proposed approach based on characteristics is that we can only consider finite time intervals on which classical solutions of the HJ equations exist. It would be of interest to extend the methodology to weak solutions, after the potential formation of singularities. It would also be of interest to combine our work with the work on curse of dimensionality with respect to dimension of Euclidean space, cited in Section 1. Furthermore, in practice we recommend learning the Hamiltonian flow, which underlies the method of characteristics for the HJ equation, using symplectic neural networks [27]. However, the analysis of these neural networks is not yet developed to the extent needed for our complexity analysis in this paper, which builds on the work in [54]. Extending the analysis of symplectic neural networks, and using this extension to analyze generalizations of HJ-Net as defined here, are interesting directions for future study. Finally, we note that it is of interest to iterate the learned operator. In order to do this, we would need to generalize the error estimates to the $C^{1}$ topology. This could be achieved either by interpolation between higher-order $C^{s}$ bounds of the proposed methodology for $s>1$ combined with the existing error analysis, or by using the gradient variable $p$ in the interpolation.
@@ A.1. Preliminaries in Finite Dimensions
Let $f:\mathbb{R}^{D}\to\mathbb{R}$ be a function. Following [54, Sect. 4.3], we will denote by $\mathcal{N}(f,\epsilon)$ the minimal number of hidden computation units that is required to approximate $f$ to accuracy $\epsilon$ by an arbitrary ReLU feedforward network $\Psi$, uniformly over the unit cube $[0,1]^{D}$; i.e. $\mathcal{N}(f,\epsilon)$ is the minimal integer $M$ such that there exists a ReLU neural network $\Psi$ with at most $M$ computation units [^4] The number of computation units equals the number of scalar evaluations of the activation $\sigma$ in a forward-pass, cf. [1]. and such that
@@
Note that the number of (non-trivial) hidden computation units $M$ of a neural network $\Psi:\mathbb{R}^{D}\to\mathbb{R}$ obeys the bounds $M\leq\mathrm{size}(\Psi)\leq 5DM^{4}$: The lower bound follows from the fact that each (non-trivial) computation unit has at least one non-zero connection or a bias feeding into it. To derive the upper bound, we note that any network with at most $M$ computation units can be embedded in a fully connected enveloping network (allowing skip-connections) [54, cf. Fig. 6(a)] with depth $M$, width $M$, where each hidden node is connected to all other $M^{2}-1$ hidden nodes plus the output, and where each node in the input layer is connected to all $M^{2}$ hidden units plus the output. In addition, we take into account that each computation unit and the output layer have an additional bias. The existence of this enveloping network thus implies the size bound
@@
The latter conclusion was established by Yarotsky [54, Thm. 5] (with $\gamma=1/9$). Proposition 2.1 improves on Yarotsky’s result in two important ways: firstly, it implies that the lower bound $\mathrm{size}(\Psi_{\epsilon})\geq\epsilon^{-\gamma D/r}$ holds for all $\epsilon\leq\overline{\epsilon}$, and not only along a (unspecified) sequence $\epsilon_{k}\to 0$; secondly, it shows that the constant $\overline{\epsilon}$ can be chosen independently of $D$. This generalization of Yarotsky’s result will be crucial for the extension to the infinite-dimensional case, which is the goal of this work.
To prove Proposition A.3, we will need two intermediate results, which build on [54]. We start with the following lemma providing a lower bound on the required size of a fixed neural network architecture, which is able to approximate arbitrary $f\in\mathring{F}_{D,r}$ to accuracy $\epsilon\geq 0$. This result is based on [54, Thm. 4(a)], but explicitly quantifies the dependence on the dimension $D$. This dependence was left unspecified in earlier work.
@@
The proof in [54] is based on the Vapnik-Chervonenkis (VC) dimension. We will not repeat the entire argument here, but instead discuss only the required changes in the proof of Yarotsky, which are needed to prove our extension of his result. In fact, there are only two differences between the proof our Lemma A.4, and [54, Thm. 4(a)], which we now point out: The first difference is that we consider $\mathring{F}_{D,r}$, consisting of $C^{r}(\mathbb{R}^{D})$ functions vanishing at the origin, whereas [54] considers the unit ball in $W^{r,\infty}([0,1]^{D})$. Nevertheless the proof of [54] applies to our setting in a straightforward way. To see this, we recall that the construction in [54, Proof of Thm. 4, eq. (35)], considers $f\in F_{D,r}$ of the form
@@
As shown in [54], the above construction allows one to encode arbitrary Boolean values $z_{1},\dots,z_{N^{D}}\in\{0,1\}$ by construction suitable $f\in\mathring{F}_{D,r}$; this in turn can be used to estimate a VC-dimension related to $\Psi({\,\cdot\,};\theta)$ from below, following verbatim the arguments in [54]. As no changes are required in this part of the proof, we will not repeat the details here; Instead, following the argument leading up to [54, eq. (38)], and under the assumptions of Lemma A.4, Yarotsky’s argument immediately implies the following lower bound
@@
While Lemma A.4 applies to a fixed architecture capable of approximating all $f\in\mathring{F}_{D,r}$, our goal is to instead construct a single $f\in\mathring{F}_{D,r}$ for which a similar lower complexity bound holds for arbitrary architectures. The construction of such $f$ will rely on the following lemma, based on [54, Lem. 3]:
@@
Lemma A.5 will be our main technical tool for the proof of Proposition A.3. We also recall that $\mathcal{N}(f,\epsilon)$ satisfies the following properties, [54, eq. (42)–(44)]:
@@ Appendix B Short-time existence of $C^{r}$-solutions
Theorem B.1 (Implicit Function Theorem, see e.g. [6, Section 2.2]).
@@ Appendix C Reconstruction from scattered data
The first lemma is the following special case of the Vitali covering lemma, a well-known geometric result from measure theory (see e.g. [52, Lemma 1.2]):
@@
Given $Q=\{q^{1},\dots,q^{N}\}$, the proof of Lemma C.1 (see e.g. [52, Lemma 1.2]) shows that the subset $Q^{\prime}\subset Q$ of Lemma C.1 can be found by the following greedy algorithm, which proceeds by iteratively adding elements to $Q^{\prime}$ (the following is in fact the basis for Algorithm 1 in the main text):
@@
This is immediate from [53, Corollary 4.8]. ∎
@@ D.1. Approximation of Hamiltonian flow
(Proposition 5.3) It follows from [54, Theorem 1] (and a simple scaling argument), that there exists a constant $C=C(r,d)>0$, such that for any $\epsilon>0$, there exists a neural network $\Psi_{t}\approx\Psi_{t}^{\dagger}$ satisfying the bound,

### v4

@@ 5.1. HJ-Net Beats The Curse of Parametric Complexity
Remark 5.2 (Non-uniform sampling).
In practice, an input-function dependent, non-uniform sampling could improve the results of the moving least squares interpolant. This is especially relevant given the potentially rapid growth of second and higher-order derivatives along characteristics. However, this would likely not lead to faster asymptotic convergence rates, motivating us to restrict attention to the uniformly sampled setting for our analysis, for simplicity.
@@
The error and complexity estimate of Theorem 5.1 implies that for moderate dimensions $d$, and for sufficiently smooth input functions $u_{0}\in\mathcal{F}\subset C^{r}_{\mathrm{per}}$, with $r>3d+1$, the overall complexity of this approach scales at most linearly in $\epsilon^{-1}$: Indeed, the mapping of the data points $(q^{j}_{0},p^{j}_{0},z^{j}_{0})\mapsto(q^{j}_{t},p^{j}_{t},z^{j}_{t})$ for $j=1,\dots,N$ requires $N$ forward-passes of the neural network $\Psi_{t}$, which has $O(\epsilon^{-(2d+1)/r}\log(\epsilon^{-1}))$ internal degrees of freedom. Since $N=O(\epsilon^{-d/r})$ the computational complexity of this mapping is thus bounded by $O(\epsilon^{-(3d+1)/r}\log(\epsilon^{-1}))=O(\epsilon^{-1})$. A naive implementation of the pruning algorithm requires at most $O(N^{2})=O(\epsilon^{-2d/r})=O(\epsilon^{-1})$ comparisons. The computational complexity of the reconstruction by the moving least squares method with $m\leq N$ (pruned) interpolation points and with $M\sim\epsilon^{-d/r}$ evaluation points can be bounded by $O(m+M)=O(\epsilon^{-d/r}+M)=O(\epsilon^{-1})$ [66, last paragraph of Chapter 4.2]. In particular, the overall complexity to obtain an $\epsilon$-approximation of the output function $\mathcal{S}_{t}(u_{0})\approx\mathcal{S}_{t}^{\dagger}(u_{0})$ at e.g. the encoding points $Q_{\epsilon}$ (with $M=|Q_{\epsilon}|\sim\epsilon^{-d/r}$) is at most linear in $\epsilon^{-1}$.
@@ 5.3. Moving Least Squares Reconstruction Error
the pruned moving least squares approximant $f_{{z},{Q}}:=f_{z,Q,n=r-1}$ of Algorithm 2 with interpolation data points $Q=\{q^{1},\dots,q^{N}\}$ and data values $z=\{z^{1},\dots,z^{N}\}$, satisfies
@@ 6. Conclusions
The first contribution of this work is to study the curse of dimensionality in the context of operator learning, here interpreted in terms of the infinite dimensional nature of the input space. We state a theorem which, for the first time, establishes a general form of the curse of parametric complexity, a natural scaling limit of the curse of dimensionality in high-dimensional approximation, characterized by lower bounds which are exponential in the desired error. The theorem demonstrates that in general it is not possible to obtain complexity estimates, for the size of the approximating neural operator, that grow algebraically with inverse error unless specific structure in the underlying operator is leveraged; in particular, we prove that this additional structure has to go beyond $C^{r}$- or Lipschitz-regularity. This considerably generalizes and strengthens earlier work on the curse of parametric complexity in [41], where a mild form of this curse had been identified for PCA-Net. As shown in the present work, our result applies to many proposed operator learning architectures including PCA-Net, the FNO, and DeepONet and recent nonlinear extensions thereof. The lower complexity bound in this work is obtained for neural operator architectures based on standard feedforward ReLU neural networks, and could likely be extended to feedforward architectures employing piecewise polynomial activations. It is of note that algebraic complexity bounds, which seemingly overcome the curse of parametric complexity of the present work, have recently been derived for the approximation of Lipschitz operators [62]; these results build on non-standard neural network architectures with either superexpressive activation functions, or non-standard connectivity, and therefore do not contradict our results. In fact, a recent follow-up [42] to the present work sheds further light on this question from an information-theoretic perspective, resulting in a statement of the curse of parametric complexity which is independent of the activation function.
@@
The second contribution of this paper is to present an operator learning framework for Hamilton-Jacobi equations, and to provide a complexity analysis demonstrating that the methodology is able to tame the curse of parametric complexity for these PDE. We present the ideas in a simple setting, and there are numerous avenues for future investigation. For example, as pointed out in subsection 3.3, one main limitation of the proposed approach based on characteristics is that we can only consider finite time intervals on which classical solutions of the HJ equations exist. As already pointed out there, it would be of interest to extend the methodology to viscosity solutions, after the potential formation of singularities. It would also be of interest to combine our work with the work on curse of dimensionality with respect to dimension of Euclidean space, cited in Section 1. Furthermore, in practice we recommend learning the Hamiltonian flow, which underlies the method of characteristics for the HJ equation, using symplectic neural networks [35]. However, the analysis of these neural networks is not yet developed to the extent needed for our complexity analysis in this paper, which builds on the work in [67]. Extending the analysis of symplectic neural networks, and using this extension to analyze generalizations of HJ-Net as defined here, are interesting directions for future study. Finally, we note that it is of interest to iterate the learned operator. In order to do this, we would need to generalize the error estimates to the $C^{1}$ topology. This could be achieved either by interpolation between higher-order $C^{s}$ bounds of the proposed methodology for $s>1$ combined with the existing error analysis, or by using the gradient variable $p$ in the interpolation.
@@ A.1. Preliminaries in Finite Dimensions
Let $f:\mathbb{R}^{D}\to\mathbb{R}$ be a function. Following [67, Sect. 4.3], we will denote by $\mathcal{N}(f,\epsilon)$ the minimal number of hidden computation units that is required to approximate $f$ to accuracy $\epsilon$ by an arbitrary ReLU feedforward network $\Psi$, uniformly over the unit cube $[0,1]^{D}$; i.e. $\mathcal{N}(f,\epsilon)$ is the minimal integer $M$ such that there exists a ReLU neural network $\Psi$ with at most $M$ computation units [^4] The number of computation units equals the number of scalar evaluations of the activation $\sigma$ in a forward-pass, cf. [3]. and such that
@@
Note that the number of (non-trivial) hidden computation units $M$ of a neural network $\Psi:\mathbb{R}^{D}\to\mathbb{R}$ obeys the bounds $M\leq\mathrm{size}(\Psi)\leq 5DM^{4}$: The lower bound follows from the fact that each (non-trivial) computation unit has at least one non-zero connection or a bias feeding into it. To derive the upper bound, we note that any network with at most $M$ computation units can be embedded in a fully connected enveloping network (allowing skip-connections) [67, cf. Fig. 6(a)] with depth $M$, width $M$, where each hidden node is connected to all other $M^{2}-1$ hidden nodes plus the output, and where each node in the input layer is connected to all $M^{2}$ hidden units plus the output. In addition, we take into account that each computation unit and the output layer have an additional bias. The existence of this enveloping network thus implies the size bound
@@
The latter conclusion was established by Yarotsky [67, Thm. 5] (with $\gamma=1/9$). Proposition 2.1 improves on Yarotsky’s result in two important ways: firstly, it implies that the lower bound $\mathrm{size}(\Psi_{\epsilon})\geq\epsilon^{-\gamma D/r}$ holds for all $\epsilon\leq\overline{\epsilon}$, and not only along a (unspecified) sequence $\epsilon_{k}\to 0$; secondly, it shows that the constant $\overline{\epsilon}$ can be chosen independently of $D$. This generalization of Yarotsky’s result will be crucial for the extension to the infinite-dimensional case, which is the goal of this work.
To prove Proposition A.3, we will need two intermediate results, which build on [67]. We start with the following lemma providing a lower bound on the required size of a fixed neural network architecture, which is able to approximate arbitrary $f\in\mathring{F}_{D,r}$ to accuracy $\epsilon\geq 0$. This result is based on [67, Thm. 4(a)], but explicitly quantifies the dependence on the dimension $D$. This dependence was left unspecified in earlier work.
@@
The proof in [67] is based on the Vapnik-Chervonenkis (VC) dimension. We will not repeat the entire argument here, but instead discuss only the required changes in the proof of Yarotsky, which are needed to prove our extension of his result. In fact, there are only two differences between the proof our Lemma A.4, and [67, Thm. 4(a)], which we now point out: The first difference is that we consider $\mathring{F}_{D,r}$, consisting of $C^{r}(\mathbb{R}^{D})$ functions vanishing at the origin, whereas [67] considers the unit ball in $W^{r,\infty}([0,1]^{D})$. Nevertheless the proof of [67] applies to our setting in a straightforward way. To see this, we recall that the construction in [67, Proof of Thm. 4, eq. (35)], considers $f\in F_{D,r}$ of the form
@@
As shown in [67], the above construction allows one to encode arbitrary Boolean values $z_{1},\dots,z_{N^{D}}\in\{0,1\}$ by construction suitable $f\in\mathring{F}_{D,r}$; this in turn can be used to estimate a VC-dimension related to $\Psi({\,\cdot\,};\theta)$ from below, following verbatim the arguments in [67]. As no changes are required in this part of the proof, we will not repeat the details here; Instead, following the argument leading up to [67, eq. (38)], and under the assumptions of Lemma A.4, Yarotsky’s argument immediately implies the following lower bound
@@
While Lemma A.4 applies to a fixed architecture capable of approximating all $f\in\mathring{F}_{D,r}$, our goal is to instead construct a single $f\in\mathring{F}_{D,r}$ for which a similar lower complexity bound holds for arbitrary architectures. The construction of such $f$ will rely on the following lemma, based on [67, Lem. 3]:
@@
Lemma A.5 will be our main technical tool for the proof of Proposition A.3. We also recall that $\mathcal{N}(f,\epsilon)$ satisfies the following properties, [67, eq. (42)–(44)]:
@@ Appendix B Short-time existence of $C^{r}$-solutions
Theorem B.1 (Implicit Function Theorem, see e.g. [11, Section 2.2]).
@@ Appendix C Reconstruction from scattered data
The first lemma is the following special case of the Vitali covering lemma, a well-known geometric result from measure theory (see e.g. [65, Lemma 1.2]):
@@
Given $Q=\{q^{1},\dots,q^{N}\}$, the proof of Lemma C.1 (see e.g. [65, Lemma 1.2]) shows that the subset $Q^{\prime}\subset Q$ of Lemma C.1 can be found by the following greedy algorithm, which proceeds by iteratively adding elements to $Q^{\prime}$ (the following is in fact the basis for Algorithm 1 in the main text):
@@
This is immediate from [66, Corollary 4.8]. ∎
@@ D.1. Approximation of Hamiltonian flow
(Proposition 5.4) It follows from [67, Theorem 1] (and a simple scaling argument), that there exists a constant $C=C(r,d)>0$, such that for any $\epsilon>0$, there exists a neural network $\Psi_{t}\approx\Psi_{t}^{\dagger}$ satisfying the bound,